OpenStudy (anonymous):

Find the maturity value. The principal is $1400 The rate is 16% Compounded is Quarterly Time is 6 years

OpenStudy (anonymous):

A=1400(1+(0.16/4))^(4*6) A=1400*1.04^24 A=1400*2.563304165 A=$3.588.63